### v1

# Understanding the Application of Utility Theory in Robotics and Artificial Intelligence: A Survey

Abstract — As a unifying concept in economics, game theory, and operations research, even in the Robotics and AI field, the utility is used to evaluate the level of individual needs, preferences, and interests. Especially for decision-making and learning in multi-agent/robot systems (MAS/MRS), a suitable utility model can guide agents in choosing reasonable strategies to achieve their current needs and learning to cooperate and organize their behaviors, optimizing the system’s utility, building stable and reliable relationships, and guaranteeing each group member’s sustainable development, similar to the human society. Although these systems’ complex, large-scale, and long-term behaviors are strongly determined by the fundamental characteristics of the underlying relationships, there has been less discussion on the theoretical aspects of mechanisms and the fields of applications in Robotics and AI. This paper introduces a utility-orient needs paradigm to describe and evaluate inter and outer relationships among agents’ interactions. Then, we survey existing literature in relevant fields to support it and propose several promising research directions along with some open problems deemed necessary for further investigations.
I INTRODUCTION
When people study, analyze, and design robotic or artificial intelligence (AI) systems, they would like to compare them with natural systems working. For example, many natural systems (e.g., brains, immune system, ecology, societies) are characterized by apparently complex behaviors that emerge as a result of often nonlinear spatiotemporal interactions among a large number of component systems at different levels of the organization [1]. From the single-robot perspective, cognitive developmental robotics (CDR) aims to provide a new understanding of how human higher cognitive functions develop by using a synthetic approach that developmentally constructs cognitive functions [2]. For multi-agent systems (MAS) [^1] In this paper, we use the term agent to mean autonomous intelligent entities, such as robot., they potentially share the properties of swarm intelligence [3] in practical applications (like search, rescue, mining, map construction, exploration, etc.), representing the collective behavior of distributed and self-organized systems. There are a lot of research fields related to it, especially building so-called artificial social systems, such as drones and self-driving cars. Specifically, in multi-robot systems (MRS), each robot can internally estimate the utility of executing an action. And robots’ utility estimates will be inexact for several reasons, including mission cost, strategy rewards, sensor noise, general uncertainty, and environmental change [4, 5].
Figure 1: Rewards (Cumulative Utilities) Driven Behaviors
Furthermore, in MAS cooperation, trust is an essential component for describing the agent’s purposes. In other words, there are complex relationships between utility, transparency, and trust, which depend on individual purposes [6]. Especially considering humans and robots collaborated as a team, recognizing the utility of robots will increase interest in group understanding and improve Human-Robot Interaction (HRI) performance in tasks [7]. In this paper, we introduce a utility-orient needs paradigm and survey existing work on robotics and AI through the lens of utility. We organize this review by the relationships between single-agent, multi-agent, trust among agents, and HRI to underscore the effects of utility on interaction outcomes.
I-A Scope and Contributions
Figure 2: Illustration of the utility-orient needs paradigm based on agents’ motivations among single-agent, MAS, and HRI.
One of the significant challenges in designing robotic and AI systems is how to appropriately describe the relationships of their interconnected elements and natures. It is natural to introduce a unifying concept expressing various components and evaluating diverse factors on the same ground. As a results, Utility Theory is a wide cross-area of research that lies at the intersection of economics, operations research, robotics, AI, cognitive science, and psychology. To the best of our knowledge, this survey is the first to formalize a comprehensive model of the utility paradigm in agents’ interaction from agents’ needs perspective and situate prior and ongoing research regarding how the utility can influence inter and outer relationships among agents and affect systems’ performance. The main contributions of this paper are listed as follow:
  • We introduce a utility paradigm based on agents’ needs to describe and evaluate inter and outer relationships among agents’ interaction, integrating and summarizing insights from robotics and AI studies.
  • We review existing literature in relevant fields that support the proposed utility-orient needs paradigm of agents’ interaction.
  • We propose several open problems for future research based on utility theory, which will support researchers in building robust, reliable, and efficient robotic and AI systems, especially in the MAS and HRI domain.
Fig. 2 outlines the vital relationships of the utility-orient needs paradigm based on agents’ motivations among single-agent, MAS, and HRI, which affect systems’ abilities to meet the requirements of specific tasks and adapt to human needs and uncertain environments. We survey how the utility-orient relationships affect the performance of different systems, especially evaluating the trust level among agents through various types of utilities and integrating human needs and agents’ utilities to build a harmonious team.
II Synopsis of Affordance Formalisms
This section summarises the evolution of formalisms that use the concept of utilities to improve agents’ performance. This evolution can be classified into two main stages. Mathematical conceptualizations characterize the first stage as extensions of theories from various fields. The second stage corresponds to formalisms that focus on optimizing systems’ capabilities by unifying different concepts based on Utility Theory.
II-A Utility Theory
Utility theory is postulated in economics to explain the behavior of individuals based on the premise that individuals can consistently rank the order of their choices depending on their needs or preferences. This theoretical approach aims to quantify an agent’s degree of preference across a set of available alternatives and understand how these preferences change when an agent faces uncertainty about which alternative it will receive [8].
To describe the interactions between multiple utility-theoretic agents, we use the specific utility function to analyze their preferences and rational action. The utility function is a mapping from states of the world to real numbers, which are interpreted as measures of an agent’s level of happiness (needs) in the given states. If the agent is uncertain about its current state, the utility is defined as the expected value (Eq. (1)) of its utility function for the appropriate probability distribution over the states [8]. From the perspective of the connection between a decision-maker and its preference, a decision-maker would rather implement a more preferred alternative (act, course of action, strategy) than one that is less preferred [9].
$\begin{split}\mathbb{E}\left[u(X)\right]=\sum_{i}u(x_{i})\cdot P(x_{i})\end{split}$ (1)
Generally, utility serves as a unified concept in value systems to describe the capacity of action, behavior, and strategy to satisfy the agents’ desires or needs in their decision-making. Furthermore, modern utility theory has its origins in the theory of expected value, and the value of any course of action could be determined by multiplying the gain realized from that action by the likelihood of receiving that gain [10].
II-B Basic Concepts in Agent’s Motivation
Supposing AI agents can sense various environments, monitor their behaviors, and get internal and external information through perception $P(t)$ in the given time. And the data can generally be categorized into two classes: different variables in the environments (like light, sound, color, etc.) and agents’ physical structure and operation (such as arm extension, contact in leg, etc), the operation of the cognitive processes within the robot (learning error, perceptual novelty, etc.) [11]. Then, the set of values transmits to the agent’s actuators making up the actuation vector $A(t)$ in a given time. Furthermore, we can define a series of specific concepts as follow:
  • Goal: An agent obtains utilities in a perceptual state $P(t)$.
  • Innate Value/needs/Motivation/Drive: An agent has the plasticity to encode the feedback received from its interactive environments continuously, and the encoding of experience should facilitate the emergence of adaptive behaviors [12].
  • Strategy: A strategy describes the general plan of an agent achieving short-term or long-term goals under uncertainty, which involves setting sub-goals and priorities, determining action sequences to fulfill the tasks, and mobilizing resources to execute the actions [13].
  • Utility: The benefit of using a strategy is to achieve a goal satisfying a specific innate value and needs. The utility is determined by the system’s interaction with the environment and is not known a priori, traditionally called Reward in the reinforcement learning field [11].
  • Expected Utility ($e_{u}$): The probability of obtaining utility starts from a given perceptual state $P(t)$, which is modulated by the amount of utility obtained.
  • Utility Model: A function provides the expected utility for any point in state space, which establishes the relationship between the real utility obtained from the interaction and how to achieve the goal.
  • Value Function (VF): A specific utility model expresses the expected utility for any perceptual state $P(t)$:
    $\begin{split}e_{u}(t+1)=VF\left(P(t+1)\right)\end{split}$ (2)
  • Episode: It describes a circle of the interaction between agents and environments:
    $\begin{split}episode=\left\{P(t);A(t);P(t+1;e_{u}(t+1)\right\}\end{split}$ (3)
Figure 3: The illustration of the model for a generic intrinsically motivated agent.
Especially, [14] presents a model for a generic intrinsically motivated agent shown in Fig. 3  [^2] Here, intrinsic motivation components modulate behavioral components to connect perception components (sensors) and actuators.. Through sensors generating an environment state at each time step, the agent will form various goals and corresponding expected utilities in motivation components. Then, the behavioral components will make decisions based on specific goals to modify the agent’s internal state and generate diverse strategies and actions to achieve goals by triggering actuators to interact with the environment changing the external state.
II-C Needs-driven Behaviors
In nature, from cell to human, all intelligent agents represent different kinds of hierarchical needs such as the low-level physiological needs (food and water) in microbe and animal; the high-level needs self-actualization (creative activities) in human being [15]. Different levels of needs stimulate all agents to adopt diverse behaviors achieving certain goals or purposes and satisfying their various needs in the natural world [16]. They also can be regarded as self-organizing behaviors [17], which keep on interacting with the environments and exchanging information and energy to support the system’s survival and development.
Specifically, the needs describe the necessities for a self-organizing system to survive and evolve, which arouses an agent to action toward a goal, giving purpose and direction to behavior and strategy. The dominant approach to modeling an agent’s interests or needs is Utility Theory. [18] introduced the self-sufficient robots’ decision-making based on a basic work cycle – find fuel – refuel. They utilized some utility criterion needs, such as utility maximization, to guide robots performing opportunistic behaviors [19].
From the utility theory perspective, needs are regarded as innate values driving agents to interact with environments. Especially the ultimate payoff for animals is to maximize inclusive fitness utilizing some evolutionary strategy, and for humans is to maximize profitability through some marketing strategy [20, 21]. If we consider the ecological status of AI agents, like robots, accomplishing tasks in the real world, their situation is similar to our ecological situation [22].
Especially reinforcement learning (RL) is an excellent model to describe the needs-driven behaviors of AI agents. The essence of RL is learning from interaction based on reward-driven (such as utilities and needs) behaviors, much like natural agents. When an RL agent interacts with the environment, it can observe the consequence of its actions and learn to change its behaviors based on the corresponding rewards received. The theoretical foundation of RL is the paradigm of trial-and-error learning rooted in behaviorist psychology [23]. In RL, the policy dictates the actions that the agent takes as a function of its state and the environment, and the goal of the agent is to learn a policy maximizing the expected cumulative rewards/utilities in the process. Fig. 1 illustrates the RL process of rewards-driven behaviors [24].
II-D Utility-orient Needs Paradigm Systems
Fig. 2 represents the relationships among single-agent, multi-agent, and humans based on the paradigm of the utility-orient needs. More specially, before an individual agent cooperates with group members, it needs to fit some basic needs, such as having enough energy, guaranteeing other agents’ safety, etc. Then it requires corresponding capabilities to collaborate with other agents, satisfy task requirements, and adapt to dynamically changing environments (Fig. 2 shows four categories of robot, carrier, supplier, observer, and executor, working as a team in the middle graph.). Especially trust among agents can be evaluated based on their current status and needs, performance, knowledge, experience, motivations, etc.
Furthermore, when AI agents work with humans, they should satisfy humans’ needs and assist humans in fulfilling their missions efficiently. Therefore, it is necessary to build a unified utility paradigm describing the needs of AI agents and humans in the common ground for evaluating their relationships like trust, which is also the pre-condition of safety, reliability, stability, and sustainability in cooperation. In the following sections, we will review existing literature from the perspective of single-agent systems, multi-agent systems (MAS), trust among agents, and human-robot interaction (HRI) to support the proposed utility-orient needs paradigm systems.
III Single-Agent Systems
For single-agent systems, cognitive robotics is the most typical research area for implementing Utility Theory. It studies the mechanisms, architectures, and constraints that allow lifelong and open-ended improvement of perceptual, reasoning, planning, social, knowledge acquisition, and decision-making skills in embodied machines [12]. Especially building a value  [^3] In this section, we use the terms value and utility interchangeably. system to mimic the “brain” of an AI agent mapping behavioral responses for sensed external phenomena is the core component of cognitive robotics, which is also an emerging and specialized sub-field in neurorobotics, robotics, and artificial cognitive systems research.
Here, the value measures an agent’s effort to expend to obtain a reward or avoid punishment. It is not hard-wired for an AI agent, even a biological entity, and the specific value system achieved through experience reflects an agent’s subjective evaluation of the sensory space. And the value mechanisms usually have been defined as the expected values, particularly in uncertain environments. Moreover, the innate value reflects an agent’s subjective evaluation of the sensory space, but the acquired value is shaped through experience during its development.
[25] defines the artificial value system for the autonomous robot as a mechanism that can reflect the effect of their experience in future behaviors. In the life of an agent, the agent starts from inherent values, and its acquired values are modified through interaction with various environments and knowledge accumulation in learning. So the acquired value is thus activity-dependent and allows the value system to become sensitive to stimuli that are not able to trigger a value-related response by themselves [26].
According to the development of cognitive robotics, the existing value systems can mainly be classified into three categories. Neuroanatomical Systems discuss the explainable biologically inspired value systems design from neuroanatomy and physiology perspectives [27]; Neural Networks Systems build more abstract models through mathematical approaches to mimic the agent’s value systems; Motivational Systems consider the model that agents interact with environments to satisfy their innate values, and the typical mechanism is reinforcement learning (RL). We review the three categories as follow:
Figure 4: The illustration of the functions of the brain.
III-A Neuroanatomical Systems
In the studies of the theory of primate nervous systems in the brain, researchers usually used robotics systems as a test-bed [28], figuring out how neurons influence each other, how long neurons activate, and the brain regions are affected [29]. Fig. 4 illustrates the functions of the brain from anatomical and physiological perspectives. Following the goals of synthetic neural modeling, they aim to forge links between crucial anatomical and physiological properties of nervous systems and their overall function during autonomous behavior in an environment [12].
For example, [27] designed a value system in the NOMAD mobile robot, driven by the object – “taste” to modulate changes in connections between visual and motor neurons, thus linking specific visual responses to appropriate motor outputs. [30] described the robotic architecture embedding of a high-level model of the basal ganglia and related nuclei based on varying motivational and sensory inputs, such as fear and hunger, to generate coherent sequences of robot behavior.
Especially based on the principles of neuromodulation in the mammalian brain, [31] presents a strategy for controlling autonomous robots. They used a cognitive robot – CARL-1, to test the hypothesis that neuromodulatory activity can shape learning, drive attention, and select actions. The experiments demonstrated that the robot could learn to approach and move away from stimuli by predicting the positive or negative value.
Recently, [32] shows the putative functions ascribed to dopamine (DA) emerging from the combination of a standard computational mechanism coupled with differential sensitivity to the presence of DA across the striatum through testing on the simulated humanoid robot iCub interacting with a mechatronic board. However, it is worth mentioning that the current trend tries to blur the line between designing value systems that follow the neuroanatomy and biology of the brain and artificial neural networks (ANN) that are biologically inspired but less anatomically accurate [12].
Figure 5: The illustration of a two-layered feedforward neural network.
III-B Artificial Neural Networks Systems
Artificial Neural Networks (ANN) are an information manager model inspired by the biological nervous systems function of the human brain. Similarly, Fig. 4 demonstrates the connection of different function modules within the human brain working as a neural network to perform various reasoning activities. Like the human brain adjusting the synaptic relationships between and among neurons in learning, ANN also needs to modify the parameters of “nodes” (neuron) to adapt to the specific scenario through a learning process. Fig. 5 illustrates a two-layered feedforward neural network.
There are several typical ANN models, such as restricted Boltzmann machines (RBMs) [33], spiking neural networks [34], adaptive resonance theory (ART) networks [35], autoencoder (AE) [36], convolutional neural networks (CNNs) [37], growing neural gases (GNGs) [38], ect. In the ANN, Nodes working as neurons are non-linear processing units receiving various signals and performing different processing steps, as depicted graphically in Fig. 5. Moreover, the AI agent’s sensor states represent its properties for the input nodes, and the output nodes express corresponding actions and behaviors in the current status. The agent needs to learn a function that can describe its future expected reward (utility) based on the specific action or strategy in the given state. Due to the reward inverse proportion to learning error, the agent is more motivated to learn the actions with higher learning error until it gets the skill. Then, it will switch to another to learn based on its needs.
Especially [39] proposes a neural network architecture for integrating different value systems with reinforcement learning, presenting an empirical evaluation and comparison of four value systems for motivating exploration by a Lego Mindstorms NXT robot. [40] introduces a model of hierarchical curiosity loops for an autonomous active learning agent by selecting the optimal action that maximizes the agent’s learning of sensory-motor correlations. From the information theory perspective, [41, 42] presents a method linking information theoretic quantities on the behavioral level (sensor values) to explicit dynamical rules on the internal level (synaptic weights) in a systematic way. They studied an intrinsic motivation system for behavioral self-exploration based on the maximization of the predictive information using a range of real robots, such as the humanoid and Stumpy robots.
[43] recently proposed the value system of a developmental robot using the RBM as the data structure. They use simulation to demonstrate the mechanism allowing the agent to accumulate knowledge in an organized sequence with gradually increasing complexity while hardly learning from purely random areas. From another angle, RL gradually becomes a popular and effective model for non-specific reward tasks [12], especially combined with deep learning methods.
III-C Motivational Systems
To clarify agents’ motivation, providing an expected utility for each perceptual state in the motivation evaluation is the pre-condition in their decision-making, especially in discovering goals achieving them automatically, and determining the priority of drives [11]. Many machine learning algorithms share the properties of value systems and provide a way for AI agents to improve their performance in tasks and respect their innate values [12]. The most typical and general model is reinforcement learning (RL).
The essence of RL is learning from interaction. When an RL agent interacts with the environment, it can observe the consequence of its actions and learn to change its behaviors based on the corresponding rewards received. Moreover, the theoretical foundation of RL is the paradigm of trial-and-error learning rooted in behaviorist psychology [23]. In this process, the best action sequences of the agent is determined by the rewards through interaction and the goal of the agent is to learn a policy $\pi$ maximizing its expected return (cumulative utilities with discount). In a word, given a state, a policy returns an action performance and learn an optimal policy from the consequence of actions through trial and error to maximize the expected return in the environment. Figure 1 illustrates this perception-action-learning loop.
Figure 6: The illustration of Agent Needs Hierarchy as five different modules.
Furthermore, the flexibility of RL permits it as an effective utility system to model agents’ motivation through designing task-nonspecific reward signals derived from agents’ experiences [39]. [44] introduced one of the earliest self-motivated value systems integrated with RL, demonstrating that the SAIL robot could learn to pay attention to salient visual stimuli while neglecting unimportant input. To understand the origins of reward and affective systems, [45] built artificial agents sharing the same intrinsic constraints as natural agents: self-preservation and self-reproduction. [46] uses a table-based Q-learning as a baseline compared with other function approximation techniques in motivated reinforcement learning.
From the mechanism of intelligent adaptive curiosity perspective, [47] introduced an intrinsic motivation system pushing a robot toward situations to maximize its learning progress. [48] presented a curiosity-driven reinforcement learning approach that explores the iCub’s state-action space through information gain maximization, learning a world model from experience, and controlling the actual iCub hardware in real time. To help the iCub robot intrinsically motivated to acquire, store and reuse skills, [49] introduced Continual Curiosity driven Skill Acquisition (CCSA), a set of compact low-dimensional representations of the streams of high-dimensional visual information learned through incremental slow feature analysis.
Especially [50] proposed an RL model combined with the agent basic needs satisfaction and intrinsic motivations. It was tested in a simulated survival domain, where a robot was engaged in survival tasks such as finding food or water while avoiding dangerous situations. Furthermore, [51, 52] recently proposed the first formalized utility paradigm – Agent Needs Hierarchy – for the needs of an AI agent to describe its motivations from the system theory and psychological perspective. The Agent Needs Hierarchy is similar to Maslow’s human needs pyramid [15]. An agent’s abstract needs for a given task are prioritized and distributed into multiple levels, each of them preconditioned on their lower levels. At each level, it expresses the needs as an expectation over the corresponding factors/features’ distribution to the specific group [52]. Fig. 6 illustrates the Agent Needs Hierarchy as five different modules in the proposed Self-Adaptive Swarm Systems (SASS) [53].
Specifically, the lowest (first) level is the safety features of the agent (e.g., features such as collision detection, fault detection, etc., that assure safety to the agent, human, and other friendly agents in the environment). The safety needs can be calculated through its safety feature’s value and corresponding safety feature’s probability based on the current state of the agent. After satisfying safety needs, the agent considers its basic needs, which includes features such as energy levels, data communication levels that help maintain the basic operations of that agent. Only after fitting the safety and basic needs, an agent can consider its capability needs, which are composed of features such as its health level, computing (e.g., storage, performance), physical functionalities (e.g., resources, manipulation), etc.
At the next higher level, the agent can identify its teaming needs that accounts the contributions of this agent to its team through several factors (e.g., heterogeneity, trust, actions) that the team needs so that they can form a reliable and robust team to successfully perform a given mission.
Ultimately, at the highest level, the agent learns skills or features to improve its capabilities and performance in achieving a given task through RL. Later work by [54] presented a strategy-oriented Bayesian soft actor-critic (BSAC) Model based on the agent needs, which integrates an agent strategy composition approach termed Bayesian Strategy Network (BSN) with the soft actor-critic (SAC) method to achieve efficient deep reinforcement learning (DRL).
Generally, the set of agent needs reflects its innate values and motivations. It is described as the union of needs at all the levels in the needs hierarchy. Each category needs level is combined with various similar needs (expected utilities) presenting as a set, consisting of individual hierarchical and compound needs matrix [55]. It drives the agent to develop various strategies and diverse skills to satisfy its intrinsic needs and values, much like a human does.
IV Multi-Agent Systems
A multi-agent system (MAS) consists of several agents interacting with each other, which may act cooperatively, competitively, or exhibit a mixture of these behaviors [56]. In the most general case, agents will act on behalf of users with different goals and motivations to cooperate, coordinate, and negotiate with each other, achieving successful interaction, such as in human society [57]. Most MAS implementations aim to optimize the system’s policies with respect to individual needs and intrinsic values, even though many real-world problems are inherently multi-objective [58]. Therefore, many conflicts and complex trade-offs in the MAS need to be managed, and compromises among agents should be based on the utility mapping the innate values of a compromise solution – how to measure and what to optimize [59].
However, in the MAS setting, the situations will become much more complex when we consider individual utility reflects its own needs and preferences. For example, although we assume each group member receives the same team rewards in fully cooperative MAS, the benefits received by an individual agent are usually significant differences according to its contributions and innate values in real-world scenarios or general multi-agent settings. Especially in distributed intelligent systems, such as multi-robot systems (MRS), self-driving cars, and delivery drones, there is no single decision-maker with full information and authority. Instead, the system performance greatly depends on the decisions made by interacting entities with local information and limited communication capabilities [60]. This section reviews the works of decision-makers compromise in MAS from the game-theoretic utility cooperative decision-making perspective based on their needs and innate values. Then, we discuss the design and influence of individual utility functions to induce desirable system behaviors and optimization criteria.
IV-A Game-theoretic Utility Systems
From a game-theoretic control perspective, most of the research in game theory has been focused on single-stage games with fixed, known agent utilities [61], such as distributed control in communication [62] and task allocation [63]. Especially, recent MAS research domains focus on solving path planning problems for avoiding static or dynamical obstacles [64] and formation control [65] from the unintentional adversary perspective. For intentional adversaries, the Pursuit Domain [66, 67] primarily deals with how to guide pursuers to catch evaders [68] efficiently. Similarly, the robot soccer domain [69] deals with how one group of robots wins over another group of robots on a typical game element.
Furthermore, optimal control of MAS via game theory assumes a system-level objective is given, and the utility functions for individual agents are designed to convert a Multi-Agent System into a potential game [70]. Although game-theoretic approaches has received intensive attention in the control community, it is still a promising new approach to distributed control of MAS. Especially dynamic non-cooperative game theory using distributed optimization in MAS has been studied in [71]. In [72], the authors investigated the MAS consensus and the work in [73] provided an algorithm for large-scale MAS optimization. However, designing utility functions, learning from global goals for potential game-based optimization of control systems, and converting the local optimization problem to an original optimization problem into a potential networked game are still open challenges [74].
It is worth mentioning that [75] and [76] introduced an architectural overview from the perspectives of decomposing the problem into utility function design and building the learning component. Specifically, the utility function design concerns the satisfaction of the results of emergent behaviors from the standpoint of the system designer, and the learning component design focuses on developing distributed algorithms coordinating agents towards one such equilibrium configuration. Recently, [77, 78] solved the utility function design of the separable resource allocation problems through systematic methodology optimizing the price of anarchy. [79] discussed the learning in games of rational agents’ coverage to an equilibrium allocation by revising their choices over time in a dynamic process.
More recently, [80, 81] proposed a new model called the Game-Theoretic Utility Tree (GUT), combining the core principles of game theory and utility theory to achieve cooperative decision-making for MAS in adversarial environments. It combines with a new payoff measure based on agent needs for real-time strategy games. By calculating the Game-Theoretic Utility Computation Unit distributed at each level, the individual can decompose high-level strategies into executable low-level actions. It demonstrated the GUT in the proposed Explore Game domain and verified the effectiveness of the GUT in the real robot testbed – Robotarium [82], indicating that GUT can organize more complex relationships among MAS cooperation and help the group achieve challenging tasks with lower costs and higher winning rates. Fig 7 illustrates a MAS (UAV) working in adversarial environments with several typical scenarios.
IV-B Cooperative Decision-making and Learning
Figure 7: Illustration of A MAS (UAV) Working in Adversarial Environments with Several Typical Scenarios
Cooperative decision-making among the agents is essential to address the threats posed by intentional physical adversaries or determine tradeoffs in tactical networks [83]. Although cooperative MAS decision-making used to be studied in many separate communities, such as evolutionary computation, complex systems, game theory, graph theory, and control theory [84], these problems are either be episodic or sequential [85]. Agents’ actions or behavior are usually generated from a sequence of actions or policies, and the decision-making algorithms are evaluated based on utility-orient policy optimality, search completeness, time complexity, and space complexity.
Furthermore, existing cooperative decision-making models use Markov decision process (MDP) and its variants [86], game-theoretic methods, and swarm intelligence [84]. They mostly involve using Reinforcement Learning (RL) and Recurrent Neural Networks (RNN) to find optimal or suboptimal action sequences based on current and previous states for achieving independent or transferred learning of decision-making policies [87, 88].
Considering decision-making in the context of cooperative MAS, the learning process can be centralized or decentralized. [89] divides it into two categories: team learning and concurrent learning.
In team learning, only one learner involves the learning process and represents a set of behaviors for the group of agents, which is a simple approach for cooperative MAS learning since the standard single-agent machine learning techniques can handle it. So considering multiple learning processes improve the team’s performance, concurrent learning is the most common approach in cooperative MAS. Since concurrent learning projects the large joint team search space onto separate, smaller subset search spaces, [90] argue that it is suitable for the domains that can be decomposed into independent problems and benefit from them. Especially for individual behaviors that are relatively disjoint, it can dramatically reduce search space and computation complexity. Furthermore, breaking the learning process into smaller chunks provides more flexibility for the individual learning process using computational resources.
Figure 8: End-to-End learning of coordinated 2 vs 2 humanoid football in MuJoCo multi-agent soccer environment
However, [89] argue that the central challenge for concurrent learning is that each learner is adapting its behaviors in the context of other co-adapting learners over which it has no control, and there are three main thrusts in the area of concurrent learning: credit assignment, the dynamics of learning, and teammate modeling.
The credit assignment problem focuses on appropriately apportioning the group rewards (utilities) to the individual learner. The most typical solution is to separate the team rewards for each learner equally or the reward changing trend of individual learners are the same. The approach divvying up among each learner and receiving rewards through joint actions or strategies is usually termed global rewards. [91] argue that global reward does not scale well to increasingly difficult problems because the learners do not have sufficient feedback tailored to their own specific actions.
In contrast, if we do not divide the team rewards equally, evaluating each agent’s performance is based on individual behaviors, which means agents do not have the motivation to cooperate and tend to greedy behaviors; these methods are the so-called local reward. Through studies different credit assignment policies, [92, 93] argue that local reward leads to faster learning rates, but not necessarily to better results than global reward. Because local rewards increase the homogeneity of the learning group, he suggests that credit assignment selection relies on the desired degree of specialization. Furthermore, [94] argues that individual learning processes can be improved by combining separate local reinforcement with types of social reinforcement.
The dynamics of learning consider the impact of co-adaptation on the learning processes. Assuming agents work in dynamic changing and unstructured environments, they need to constantly track the shifting optimal behavior or strategy adapting to various situations, especially considering the agents may change other group members’ learning environments. Evolutionary game theory provides a common framework for cooperative MAS learning. [95, 96] studied the properties of cooperative coevolution and [97] visualize basins of attraction to Nash equilibria for cooperative coevolution.
Moreover, lots of research in concurrent learning involve game theory to investigate MAS problems [98, 8]. Especially for Nash equilibria, it provides a joint strategy for each group member, which lets individuals not have the motivation to shift their strategy or action in the current situation in terms of a better reward. In the full-cooperative scenario with global reward, the reward affects the benefits of each agent and only needs to consider the globally optimal Nash equilibrium. However, the relationships among agents are not clear for more general situations, which combine the cooperative and non-cooperative scenarios. In other words, individual rewards do not directly relate to the team reward, so the general-sum game [8] is the cooperative learning paradigm.
IV-C From Individual Utility to Social Welfare
As we discussed above, the reward same as the utility, reflects the agents’ diverse intrinsic motivations and needs. Individual and team rewards also directly describe the relationships between each group member and the team. We are taking MAS football and the Robot-soccer domain as an example. When considering multi-agent systems (MAS) cooperation as a team, an individual agent must first master essential skills to satisfy low-level needs (safety, basic, and capability needs) [51]. Then, it will develop effective strategies fitting middle-level needs (like teaming and collaboration) to guarantee the systems’ utilities [52]. Through learning from interaction, MAS can optimize group behaviors and present complex strategies adapting to various scenarios and achieving the highest-level needs, and fulfilling evolution. By cooperating to achieve a specific task, gaining expected utility (reward), or against the adversaries decreasing the threat, intelligent agents can benefit the entire group development or utilities and guarantee individual needs and interests. It is worth mentioning that [99] developed the end-to-end learning implemented in the MuJoCo multi-agent soccer environment [100], which combines low-level imitation learning, mid-level, and high-level reinforcement learning, using transferable representations of behavior for decision-making at different levels of abstraction. Figure 8 represents the training process.
However, suppose we extend the scalar to the entire system, like community and society, and consider optimizing the overall social welfare across all agents. In that case, we have to introduce the social choice utility. In this setting, agents have different agendas driven by their needs, leading to complex dynamic behaviors when interacting, which causes the utility hard to predict, let alone optimize. To solve this problem, we usually assume that agents are self-interested and optimize their utilities from the perspective of socially favorable desires and needs [58]. By defining the social choice function representing social welfare reflecting each agent’s needs and desired outcome, we can design a system of payments making the joint policy converge to the desired outcome [56].
For example, the government controls the parameters of tenders to balance the different requirements of projects for social welfare. Also, [101] discussed computing a convex coverage set for a cooperative multi-objective multi-agent MDP balancing traffic delays and costs as a posteriori in traffic network maintenance planning.
From another perspective, considering individual rewards and utilities from the standpoint of social choice, they are weighed and summed up, consisting of a social welfare function. In other words, the social welfare function relies on individual values or return rewards and its utility computing model or mechanism. For example, [102] discussed which properties must satisfy a multi-criteria function so it can be used to determine the winner of a multi-attribute auction. Here, social welfare may depend on attributes of the winning bids, as well as a fair outcome in terms of payments to the individual agents, that together with the costs the agents need to support to execute their bids, if chosen, typically determine the individual utilities [58].
Generally speaking, the design of the social welfare function aims to find a mechanism for making agents’ utility function transparently reflect their innate values and needs, which can optimize the joint policy concerning sustainable social welfare and guarantee individual development.
V Trust among Agents
Trust describes the interdependent relationship between agents [103], which can help us better understand the dynamics of cooperation and competition, the resolution of conflicts, and the facilitation of economic exchange [104]. From the economic angle, [105] has regarded trust as an expectation or a subjective probability and defined it using expected utility theory combined with concepts such as betrayal premium.
In Automation, authors in [106] describe trust in human-automation relationships as the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability. The trustor here is a human and the trustee usually is the automation system or an intelligent agent like robot. Those systems’ primary purpose is to assess and calibrate the trustors’ trust beliefs reflecting the automation system’s ability to achieve a specific task.
In Computing and Networking, the concept of trust involves many research fields, such as artificial intelligence (AI), human-machine interaction, and communication networks [107, 108, 109]. They regard an agent’s trust as a subjective belief which represents the reliability of the other agents’ behaviors in a particular situation with potential risks. Their decisions are based on learning from the experience to maximize its interest (or utility) and minimize risk [110]. Especially in the social IoT domain [111, 112], trust is measured based on the information accuracy and agents’ intentions (friendly or malicious) according to the current environment to avoid attacks on the system. [113] proposes a decision-making model named MATE, which applies the expected-utility decision rule to decide whether or not pervasive devices are in their node interests to load a particular component.
In MAS, trust is defined by mainly two major ways. One way is reliable relations between human operators and the MAS/MRS. In this context, trust reflects human expectations and team performance assessments [114, 115]. The primary research includes using the trust to measure the satisfactory degree of robot team performance that hybrid considering in-process behaviors and final mission accomplishments. In this process, the visible and invisible behaviors regarding humans supervising robots and robots’ behaviors will be investigated to decide the most influential factors. Especially investigating human trust in robots will help the robots to make decisions or react to emergencies, such as obstacles, when human supervision is unavailable [116, 117].
The second type of trust will be defined as reliable relations among agents, reflecting the trusting agents’ mission satisfaction, organizational group member behaviors and consensus, and task performance at the individual level. Moreover, trust has been associated with system-level analysis, such as reward and function at the organized group structure in a specific situation, the conflict solving of heterogeneous and diverse needs, and the system evolution through learning from interaction and adaptation.
On the other hand, the act of trusting by an agent can be conceptualized as consisting of two main steps: 1) trust evaluation, in which the agent assesses the trustworthiness of potential interaction partners; 2) trust-aware decision-making, in which the agent selects interaction partners based on their trust values [118]. Also, assessing the performance of agent trust models is of concern to agent trust research [119].
It is worth noticing that reputation and trust mechanisms have gradually become critical elements in MAS design since agents depend on them to evaluate the behavior of potential partners in open environments [120]. As an essential component of reputation, trust generates the reputation from agents’ previous behaviors and experiences, which becomes an implicit social control artifact [121] when agents select partners or exclude them due to social rejection (bad reputation).
Furthermore, considering the interactions between human agents and artificial intelligence agents like human-robot interaction (HRI), building stable and reliable relationships is of utmost importance in MRS cooperation, especially in adversarial environments and rescue missions [52]. In such multi-agent missions, appropriate trust in robotic collaborators is one of the leading factors influencing HRI performance [122]. In HRI, factors affecting trust are classified into three categories [123]: 1) Human-related factors (i.e., including ability-based factors, characteristics); 2) Robot-related factors (i.e., including performance-based factors and attribute-based factors); 3) Environmental factors (i.e., including team collaboration, tasking, etc.). Although there is no unified definition of trust in the literature, researchers take a utilitarian approach to defining trust for HRI adopting a trust definition that gives robots practical benefits in developing appropriate behaviors through planning and control [124, 125, 126].
Moreover, with the development of the explanation of machine learning, such as explainable AI, although explanation and trust have been intertwined for historical reasons, the utility of machine learning explanations will have more general and consequential [127]. Especially the utility of an explanation must be tied to its appropriate trust model.
Although modeling trust has been studied from various perspectives and involves many different factors, it is still challenging to develop a conclusive trust model that incorporates all of these factors. Therefore, future research into trust modeling needs to be more focused on developing general trust models based on measures other than the countless factors affecting trust [122]. Specifically, few works from the literature considered trust in socio-intelligent systems from an AI agent perspective to model trust with agents having similar needs and intrinsic values.
Recently, [128] proposed a novel trust model termed relative needs entropy (RNE) by combining utility theory and the concept of relative entropy based on the agent’s needs hierarchy [51, 52] and innate values. Similar to information entropy, it defines the entropy of needs as the difference or distance of needs distribution between agents in a specific scenario for an individual or group.
From a statistical perspective, the RNE can be regarded as calculating the similarity of high-dimensional samples from the agent needs vector. A lower RNE value means that the trust level between agents or groups is higher because their needs are well-aligned and there is low difference (distance) in their needs distributions. Similarly, a higher RNE value will mean that the needs distributions are diverse, and there exists a low trust level between the agent or groups because of their misalignment in their motivations.
As discussed above, trust is the basis for decision-making in many contexts and the motivation for maintaining long-term relationships based on cooperation and collaboration [110]. Although trust is a subjective belief, it is a rational decision based on learning from previous experience to maximize its interest (or utility) and minimize risk by choosing the best compromise between risk and possible utility from cooperation. Trust assessment involves various factors, particularly utility and risk analysis under dynamic situations, which rely on context and balance key tradeoffs between conflicting goals to maximize decision effectiveness and task performance.
VI Human-Robot Interaction
As the higher-level intelligent creature, humans represent more complex and diversified needs such as personal security, health, friendship, love, respect, recognition, and so forth. In this regard, humans are highly flexible resources due to the high degree of motion of the physical structure and the high level of intelligence and recognition [129]. When we consider humans and robots work as a team, organizing their needs and getting a common ground is a precondition for human-robot collaboration in complex missions and dynamic changing environments. Considering building more stable and reliable relationships between humans and robots, [130] depicts an updated model in HRI termed Human Robot Team Trust (HRTT). It considers the human as a resource with specific abilities that the resource can gain or lose trust with training and realizes the robots as resources with known performance attributes included as needs in the design phase of the collaborative process.
Due to the constantly increasing need for robots to interact, collaborate and assist humans, Human-Robot Interaction (HRI) poses new challenges, such as safety, autonomy, and acceptance issues [131]. From a robot needs perspective [51], it first needs to guarantee human security and health, such as avoiding collision with humans, protecting them from risks, etc. Moreover, in the higher level teaming needs, robots should consider human team members’ specialty and capability to form corresponding heterogeneous Human-Robot team adapting specific missions automatically [52]. Furthermore, efficient and reliable assistance is essential to the entire mission process in HRI. More importantly, designing an Interruption Mechanism can help humans interrupt robots’ current actions and re-organize them to fulfill specific emergency tasks or execute some crucial operations manually.
Especially humans also expect robots to provide safety and a stable working environment in missions. [132] supports the utility of people’s attitudes and anxiety towards robots to explain and predict behavior in human-robot interaction and suggests further investigation of which aspects of robots evoke what type of emotions and how this influences the overall evaluation of robots. From the psychophysiology perspective, [133] combines subjective research methods, such as behavioral and self-report measurements, to achieve a more complete and reliable knowledge of the person’s experience. For example, when people interact with AI agents like robots, which will be affected by mood and social desirability in a specific issue. However, psychophysiological techniques have different challenges in data acquisition and interpretation. Moreover, the monitor quality is influenced by both confounding environmental factors, such as noise or lightning, and by the individual psychological internal state during the evaluation, which might undermine the reliability and the correct interpretation of the gathered data [134].
From another angle, recognizing the potential utility of human affect (e.g., via facial expressions, voice, gestures, etc.) and responding to humans is essential in a human-robot team’s collaborative tasks, which would make robots more believable to human observers [135]. [136] proposed an architecture including natural language processing, higher-level deliberative functions, and implementing “joint intention theory”. [137] demonstrated that expressing affect and responding to human affect with affect expressions can significantly improve team performance in a joint human-robot task.
For Multi-agent HRI systems, they require more complex control strategies to coordinate several agents that may be dissimilar to one another (in roles, communication modes etc.), can involve interactions that connect more than two agents at once, and need special attention to address any conflicts that may arise from their interactions [138]. Specifically, the human would give high-level commands to the robot group or control them individually using a screen-based interface, and robots need to adapt to the operator’s actions or optimize some utility function (computational characteristics like task requirements, teaming needs, etc.) (See Fig. 2).
Moreover, building the shared mental models (SMM) enable team members to draw on their own well-structured common knowledge to select actions that are consistent and coordinated with those of their teammates, which are strongly correlated to team performance [139]. [140] designs the human-robot cross-training inspired by SMM to compute the robot policy aligned with human preference (needs) by iteratively switching roles between humans and robots to learn a shared plan for a collaborative task. Furthermore, considering humans infer the robot’s capabilities and partially adapt to the robot, [141] presents a game-theoretic model of human partial adaptation to the robot, where the human responds to the robot’s actions by maximizing a reward function that changes stochastically over time, capturing the evolution of their expectations of the robot’s capabilities.
In the HRI, adaptability is a crucial property provided by the SMM, which intrinsically relates to performance and is objectively measured by an adaptable controller as an independent variable to compare alongside other controllers [139]. Especially in complex, uncertain, and dynamically changing environments, robots need to adapt efficiently to humans’ behaviors and strategies caused by their frequently changing needs and motivations. Additionally, [142] showed that a robot adapting to the differences in humans’ preferences and needs is positively correlated with trust between them. Moreover, SMM improves trust and reliability between humans and robots by alleviating uncertainty in roles, responsibilities, and capabilities in the interactions. By integrating research regarding trust in automation and describing the dynamics of trust, the role of context, and the influence of display characteristics, [106] believes trust involves the experience and knowledge of humans about the needs, motivation, function processing, and performance of robots [123], such as whether robots’ capabilities meet expectations of humans and tasks [143], and minimizing system fault occurrence, system predictability, and transparency [144].
Generally speaking, through formulizing the motivations and needs of humans and robots and describing their complex, diverse, and dynamic relationships in a common ground like trust, utility theory provides a unified model to evaluate the performance of the interaction between humans and AI agents effectively and make a fundamental contribution to safety in the interaction.
VII Insights and Challenges
Currently, with the tremendous growth in AI technology, Robotics, IoT, and high-speed wireless sensor networks (like 5G), it gradually forms an artificial ecosystem termed artificial social systems that involves entire AI agents from software entities to hardware devices [53]. How to integrate artificial social systems into human society and coexist harmoniously is a critical issue for the sustainable development of human beings, such as assistive and healthcare robotics [145], social path planning and navigation [146], search and rescue [147, 52], and autonomous driving [148]. Especially the future factory is likely to utilize robots for a much broader range of tasks in a much more collaborative manner with humans, which intrinsically requires operation in proximity to humans, raising safety and efficiency issues [139].
At this point, building a robust, stable, and reliable trust network between humans and AI agents to evaluate agents’ performance and status in a common ground is the pre-condition for efficient and safe interaction in their collaboration. Here, utility is the key concept representing the individual agent’s innate values, needs and motivations over states of the world and is described as expected rewards. It measures sensitivity to the impact of actions on trust and long-term cooperation and is efficient enough to allow AI agents like robots to make real-time decisions [149]. For example, self-driving cars might be the first widespread case of trustworthy robots designed to earn trust by demonstrating how well they follow social norms.
From the learning perspective, the individual learning model can be regarded as constructing models of the other agents, which takes some portion of the observed interaction history as input, and returns a prediction of some property of interest regarding the modeled agent [150]. Taking human society as an example, human beings represent more complex and diversified needs, such as personal security, health, friendship, love, respect, recognition, and self-actualization. These diverse needs can be expressed as corresponding rewards or utility mechanisms driving people to adopt various strategies and skills to achieve their requirements. Especially the multi-layer and diverse needs of people are the foundation for describing their complex and dynamic relationships, such as trust.
Similarly, how to design a general and standard utility mechanism integrating humans and AI agents’ needs and learning to efficiently build reliable and stable trust relationships between agents and humans is essential [128]. Especially a suitable knowledge graph combined with efficient DRL can help agents learn to adopt appropriate behaviors or strategies benefiting the group utilities, adapting to human members’ needs, and guaranteeing their development in various situations [54]. Another interesting direction is behavior manipulation, known as policy elicitation, which refers to problems in human-robot collaboration wherein agents must guide humans toward an optimal policy to fulfill a task through implicit or explicit communication [151, 152]. For example, it is used to achieve effective personalized learning in AI agents’ teaching and coaching [153].
For future MAS research, AI agents learn and adapt to human needs and maintain trust and rapport among agents, which are critical for task efficiency and safety improvement [147]. And we list potential directions as follow:
  • Adopting suitable formation to perceive and survey environments predicting threats warning human team members.
  • Reasonable path planning adaptation in various scenarios avoids collision guaranteeing human security and decreasing human working environment interference.
  • Combining the specific capabilities and needs of agents and humans, calculating sensible strategies to efficiently organize the entire group collaboration and fulfill the corresponding mission.
As discussed above, Agent Needs Hierarchy is the foundation for MAS learning from interaction [55]. It surveys the system’s utility from individual needs. Balancing the needs and rewards between agents and groups for MAS through interaction and adaptation in cooperation optimizes the system’s utility and guarantees sustainable development for each group member, much like human society does. Furthermore, by designing suitable reward mechanisms and developing efficient DRL methods, MAS can effectively represent various group behaviors, skills, and strategies to adapt to uncertain, adversarial, and dynamically changing environments [53].
VIII Conclusion
This paper proposes a utility-orient needs paradigm based on agents’ needs to describe and evaluate inter and outer relationships among agents’ interactions in robotics and AI. We review existing literature illustrating the application of utility theory as a unified concept from the perspective of single-agent, MAS, agent trust, and HRI. Then, we discuss the insight and open problems for future research based on utility theory. We anticipate that this comprehensive overview of the role of the utility-orient needs paradigm based on agents’ needs in robotics and AI systems will support future research that leverages, compares, or constructs interactions for artificial social systems.